OpenStudy (anonymous):

if sec(θ)= 5/3 and the terminal point determined by θ is in quadrant 4, then CHECK ALL THAT APPLIES A. cos (θ)=3/5 B. csc (θ)= 5/4 C. tan (θ)= 4/3 D. sin (θ) = -(2/5)

OpenStudy (danjs):

sec(theta) = 1 / cos(theta)

OpenStudy (anonymous):

what does that mean

OpenStudy (danjs):

sec(theta) = 1 / cos(theta) = 5 / 3 solve for cos(theta)

OpenStudy (danjs):

in quadrant 4, cos(theta) is positive

OpenStudy (danjs):

( x, y ) = ( cos(theta) , sin(theta) ) 4th quadrant, positive X value, positive cos value

OpenStudy (anonymous):

so just A is right?

OpenStudy (danjs):

what is sin(theta 0if cos(theta) = 3/5 ?

OpenStudy (danjs):

sin^2(theta) + cos^2 (theta) = 1

OpenStudy (anonymous):

sin theta= 4/5

OpenStudy (anonymous):

Are A and B the corect answers?

OpenStudy (danjs):

remember it is the 4th quadrant, the sin will be negative

OpenStudy (anonymous):

can you just tell me if im right

OpenStudy (danjs):

you know sin and cos, check the other 2 answers B and C csc = 1 / sin = 5/4 ?

OpenStudy (danjs):

tan = sin / cos = 4 / 3 ?

OpenStudy (anonymous):

it cant be C because tangent has to be negative

OpenStudy (danjs):

yes!

OpenStudy (anonymous):

I already knew that. I am asking if A and B are the correct answers

OpenStudy (danjs):

yes, we showed cos(theta) = 3 /5 so A is right. You calculated sin(theta) = -4/5

OpenStudy (danjs):

so csc(theta) = 1 / sin(theta) = -5/4

OpenStudy (anonymous):

so are u saying B is right ?

OpenStudy (danjs):

no, csc(theta) is negative

OpenStudy (danjs):

sin is -4/5 so 1 / sin is also negative

OpenStudy (anonymous):

so csc has to be -5/4 to be true

OpenStudy (danjs):

yeah

OpenStudy (anonymous):

thank you.

OpenStudy (danjs):

so A is the only one it looks like

OpenStudy (danjs):

welcome